OpenStudy (anonymous):

When attempting to do his homework, Mark incorrectly uses the addition property of equality below. Which statement would correct Mark’s error? 1) 8 + x = 2 2) 8 – 8 + x = 2 3) x = 2 A. To balance the equation, Mark should also subtract 2 from the right side of the equation in step 2. B. To balance the equation, Mark should also subtract 2 from both sides of the equation in step 2. C. To balance the equation, Mark should also subtract 8 from the right side of the equation in step 2. D. To balance the equation, Mark should also add 8 to both sides of the equation in step 2.

OpenStudy (paki):

Basically, what you do to one side of the equation, you do to the other. 8 + x = 2 8 + x (-8) = 2 (-8) x = 2 - 8 x = -6 .... so option C works here...

OpenStudy (anonymous):

Thanks would you mind helping me with other questions?? please

OpenStudy (paki):

option A...